1) find F(0) when f(x)=x^2-5x-7
2) Find F(17) when f(x)=|x-16|


all help is appreciated

Answer by bucky(2189)

find f(0) when f(x)=x^2-5x-7
.
This problem just tells you to take the given function of x, to substitute zero for x in
that function, and then evaluate the function.
.
So, start with:
.
f(x) = x^2 - 5x - 7
.
substitute zero for x and you get:
.
f(0) = 0^2 - 5*0 - 7
.
The two terms that contain zero on the right side disappear because they each involve a
multiplication by zero, and you are left with the answer of:
.
f(0) = -7
.
Next problem:
.
Find f(17) when f(x)=|x-16|
.
Start with the function:
.
f(x) = |x - 16|
.
then substitute + 17 for every x you see to get:
.
f(17) = |17 - 16|
.
combine the two terms inside the absolute value signs to get:
.
f(17) = |+1|
.
and the absolute value of a quantity is the positive value of the quantity. So the answer
is:
.
f(17) = +1
.
Hope this helps you to understand how to evaluate functions and how to interpret what the
notation means.
.